Complete the following:

(a) 1400 ft2 = _________ yd2
(b) 1/9 yd3 = _________ ft3
(c) 4.5 lb = _________ oz
(d) 32 °C = _________ °F

Answer by checkley77(12844)

(a) 1400 ft^2 = 1,400/9=155 5/9 yd^2
(b) 1/9 yd^3 = 3 ft^3
(c) 4.5 lb = .28125 oz
(d) 32 °C = 5/9(F-32)
32C=(5F-5*32)/9
32C=(5F-160)/9
5F-160=9*32C
5F=288C+160
F=(288C+160)/5
F=288C/5+32 ANS.